OpenStudy (anonymous):

The scores received by students completing a geology exam are normally distributed. If the average score is 122 and the standard deviation is 35, what percentage of students scored between 52 and 122? (A) 34.0% (B) 37.5% (C) 42.0% (D) 47.5% (E) 48.0%

OpenStudy (anonymous):

@shinebrightlikeadimon @KyanTheDoodle

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@Hero

OpenStudy (plasmataco):

im bad at these, only 13, but I can try.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

any help is better than none

OpenStudy (plasmataco):

dunno. sry

OpenStudy (ybarrap):

$$ P(52\le X\le 122)\\ =P(52-\mu\le X-\mu\le 122-\mu)\\ =P(\cfrac{52-\mu}{\sigma}\le \cfrac{X-\mu}{\sigma}\le \cfrac{122-\mu}{\sigma})\\ $$ Does this make any sense?

OpenStudy (ybarrap):

This is converting the probability statement into a z-score, a normal random variable with mean 0 and standard deviation 1. This allows us to use the z-table to figure out the probability.

OpenStudy (ybarrap):

$$ =P(\cfrac{52-\mu}{\sigma}\le \cfrac{X-\mu}{\sigma}\le \cfrac{122-\mu}{\sigma})\\ =P(-2\le \cfrac{X-\mu}{\sigma}\le 0)\\ =P(0\ge \cfrac{X-\mu}{\sigma}\le 2)\\ $$ The last equation follows from symmetry Do you see what we did here?

OpenStudy (ybarrap):

This should be sufficient